# Law of Vision Representation in MLLMs

### v1

@@ Results.
In Table 1, we show that AC policy consistently predicts the optimal vision representation with minimal resources, given a finite search space—in this case, 13 settings. Our goal is to finetune only a small subset of the search space while still identifying the optimal vision representation within the top-3 predictions (Recall@3). However, if we randomly select a subset to train on, we need 12 out of 13 finetuning to achieve over 90% Recall@3. In contrast, the AC policy requires only 3.88 full training runs on average to reach 89.69% Recall@3. Refer to Section 5.5 for details.
5 Empirical Result Details
5.1 Experiment Settings
@@
Table 2: Vision representation explored.
For our specific MLLM pipeline, we follow the training procedure, general architecture, and dataset outlined in LLaVA (Liu et al., 2024e). The training process consists of two stages: in the first stage, we train a 2-layer GeLU-MLP connector using the LLaVA 1.5 dataset with 558K samples. In the second stage, we train both the connector and the language model on the expanded LLaVA 1.5 dataset with 665K samples. It is important to note that for each training, all factors are held constant except for the vision representation being changed. The 13 vision representation in this paper are outlined in Table 2. The MLLM benchmarks used in this paper includes 4 vision-based benchmarks, MMBench (Liu et al., 2023), MME (Fu et al., 2023), OKVQA (Marino et al., 2019), SEED-Bench (Li et al., 2024a), and 4 QCR-based benchmarks including, MMMU (Yue et al., 2024), TextVQA (Singh et al., 2019), VizWiz (Gurari et al., 2018), ScienceQA (Lu et al., 2022).
@@ 5.2 AC Score
To compute the cross-modal alignment score, we perform stage 1 training with all the vision representations to obtain the MLPs. This process requires significantly less computation than stage 2, involving only 0.298% of the trainable parameters. The alignment score for each benchmark is averaged across 100 randomly sampled images. For the correspondence score, we follow common practices using the SPair-71k (Min et al., 2019) dataset. Consequently, each benchmark has its own alignment score, while the correspondence score remains consistent across all representations.
@@ 5.3 Feature Extraction
Diffusion model is primarily used for generating images via multi-step denoising, yet a recent trend is to use diffusion model as the vision representation model (Xu et al., 2024; 2023; Zhang et al., 2024b; Tong et al., 2024a). Specifically, for diffusion models, given an image $I\in\mathbb{R}^{H\times W\times 3}$, we first add noise to the VAE-encoded representation of $I$:
@@ 5.4 Additional Results on the Law of Vision Representation
Random 2.94% 5.08%
A Score 50.80% 55.06%
C Score 46.12% 19.45%
AC Score 83.24% 66.18%
@@
Random 45.09% 35.77%
A Score 76.56% 77.45%
C Score 56.91% 29.07%
AC Score 95.72% 85.21%
@@
In Section 3, we demonstrate the strong correlation between the AC score and MLLM performance by analyzing the coefficient of determination ($R^{2}$) obtained from fitting a linear regression model. In this section, we further ablate the experiments by adding baselines, fitting model performance with random scores, A scores, and C scores separately. Additionally, we explored the relationship between the A score and C score by applying two different data transformations: no transformation and second-degree polynomial transformation. We avoid higher-degree transformations to prevent overfitting, which could obscure the true relationship between A and C scores.
As shown in Table 3, the results indicate that using the AC score consistently outperforms all other settings in terms of $R^{2}$ values. While this observation holds regardless of transformation, applying a second-degree polynomial transformation to the A and C scores yields the highest correlation with model performance. This suggests an inherent trade-off between A and C scores: vision representations with high cross-modal alignment often exhibit lower correspondence, and vice versa.
Interestingly, we observe a lower correlation between OCR-based benchmark performance and C scores, which leads to a reduced correlation between the AC score and OCR-based benchmark performance. In Section 6.3, we discuss how the use of the SPair-71k correspondence dataset across all benchmarks fails to adequately capture correspondence in images containing text.
@@ 5.5 Additional Results on the AC Policy
In Section 4, we demonstrate that fitting the AC score consistently predicts the optimal vision representation with minimal resources, given a finite search space—in this case, 13 settings. In this section, we provide detailed visualization for Table 1.
@@
When performing ablation experiments on vision encoders, it’s common to randomly select a subset to train on. However, as shown in Figure 3, with 1000 runs of simulated ablation experiments, we found that to include the optimal vision representation 81.2% of the time, at least 11 out of the 13 settings need to be trained. This suggests that running a small subset of vision representations is unreliable, especially as the search space expands, making it increasingly unlikely to identify the true optimal representation by training only a subset.
@@
In contrast, the AC policy requires only 3.88 full training runs on average to reach 89.69% Recall@3. For the most successful prediction benchmark, ScienceQA, the policy successfully identifies the optimal configuration among the top three choices in 96.6% of cases, with only three language model finetuning runs across a 13-setting search space. This result shows that AC policy significantly reduces the effort and cost of exploring vision representations for MLLMs.
6 Discussion and Limitations
6.1 Finding Vision Representations with High AC
Figure 4: Normalized performance improvements across 8 benchmarks by increasing resolution and combining features, thereby enhancing correspondence and cross-modal alignment.
Since AC score is highly correlated with MLLM benchmark performance, to improve MLLM from the vision side, it is essential to identify vision representations with high AC scores and add them into the search space. We suggest two strategies to achieve this: increasing resolution and combining features.
Figure 4 shows the normalized performance improvement summed across eight benchmarks. Increasing the resolution of well-aligned features directly enhances correspondence. For example, increasing the image resolution from 224 to 336 for CLIP, while maintaining cross-modal alignment, resulted in a performance increase from 7.1 to 7.3 out of 8.
Additionally, feature combination—merging two features with high A and C scores along the channel dimension—can enhance cross-alignment while preserving correspondence. We chose to concatenate features along the channel dimension to preserve context length and to align with the intuition from our high correspondence attention proof in Appendix A.2. Specifically, the high correspondence feature can support the retrieval of information in the attention mechanism for the feature with high cross-modal alignment. Combining CLIP with DINOv2, for instance, leads to an increase in cross-modal alignment, thereby improving model performance.
6.2 Refining A Score Design
We design the A score calculation as the maximum cosine similarity between each pair of embedding vectors from the CLIP embedding and the target embedding. However, correspondence effects can be unintentionally included if the features obtained from CLIP and the target encoder differ in resolution. For example, the A score computed for CLIP@224 with CLIP@336 is not 1, which it is supposed to be, as our assumption is that CLIP always has an A score of 1. This shows that correspondence is not fully disentangled in the A score calculation in this scenario. The best practice is to always use the same input resolution for both CLIP and the target encoder. However, since we only have CLIP at resolutions 224 and 336, we cannot compute the alignment score if the target encoder operates at a higher resolution. To mitigate this effect, we compute the averaged A score for each target vision representation using both CLIP@224 and CLIP@336, aiming to average out the influence of correspondence.
Another limitation is the use of CLIP as a reference metric. This could be problematic if another encoder with better cross-modal alignment exists in the search space. However, we believe the error should not be significant enough to cause the predicted optimal vision representation to fall outside the top-3 or top-5. If a vision representation with the same correspondence as CLIP yields significantly higher MLLM performance than CLIP, it would be advisable to switch to this model as a more accurate cross-modal alignment reference model.
6.3 Refining C Score Design
Figure 5: Visualization of correspondence on natural images and images containing text for CLIP and DINOv2. The left image shows the source image, while the right one shows the target image. Blue dots indicate key points with the highest similarity, and the green areas represent regions with relatively high similarity.
We have noted that OCR-based benchmark performance shows a weaker correlation with the AC score compared to vision-based benchmarks. The primary cause of this discrepancy lies in the correspondence dataset we selected to compute the C score. The SPair-71k dataset measures feature correspondence for natural images, such as objects like cats and trains. As shown in Figure 5, the CLIP encoder demonstrates poorer correspondence in natural images compared to DINOv2. However, when it comes to images containing text, CLIP exhibits significantly better correspondence than DINOv2 or any other encoders. Therefore, using SPair-71k to calculate the C score does not accurately capture true correspondence across all scenarios. Ideally, each benchmark should have its own keypoint-labeled images for correspondence evaluation. At a minimum, an OCR-specific correspondence dataset would be highly beneficial for assessing MLLMs. To our knowledge, no such dataset currently exists. We encourage further investigation in this direction, as it would be valuable across fields in MLLM, particularly for understanding tables and charts—a fundamental capability.
@@ A.1 Theoretical Justification of Vision Representation with High Cross-modal Alignment
Since our language model $f$ is well-trained and pre-normed, the input space to each transformer layer is bounded and compact, meaning that the values of the input are bounded by $c$, a small constant. This implies that the continuously differentiable function $f$ is Lipschitz (Kim et al., 2021). This property ensures that small changes in the input of the language model of the MLLM result in small, controlled changes in the output:
$\|f([E_{c}^{\text{image}},E_{1},E_{2},\dots,E_{n}])-f([E_{c}^{\text{text}},E_{1},E_{2},\dots,E_{n}])\|$ $\leq L\|[E_{c}^{\text{image}},E_{1},E_{2},\dots,E_{n}]$ (8)
$\quad-[E_{c}^{\text{text}},E_{1},E_{2},\dots,E_{n}]\|$ (9)
$\leq L\epsilon$ (10)
@@ A.3 All Settings Benchmark Performance
In this section, we present the performance results of all 13 vision representation settings, as summarized in Table 4. The benchmarks we evaluated include:
@@
  • SEED-Bench (Li et al., 2024a): A benchmark consisting of multiple-choice questions designed to assess both spatial and temporal understanding.
CLIP@336 CLIP@224 OpenCLIP DINOv2 SDim SD1.5 SDXL MMBench 64.26 64.18 63.406 58.51 52.84 42.53 43.73 MME 1502.70 1449.64 1460.28 1295.47 1205.33 1163.90 1212.69 MMMU 35.0 36.2 37.2 34.6 33.7 33.9 32.8 OKVQA 53.20 56.13 56.36 54.78 46.04 39.14 41.78 TextVQA 46.04 42.67 40.13 14.27 13.77 11.64 11.81 VizWiz 54.27 51.69 52.11 49.69 47.33 50.14 47.14 ScienceQA (Full) 69.97 70.23 69.78 68.17 68.66 66.75 67.72 ScienceQA 70.31 70.90 70.69 69.60 69.77 68.38 68.90 ScienceQA (Img) 69.26 68.82 67.87 65.15 66.34 63.31 65.25 SEED-Bench (Img) 66.09 65.13 64.71 61.39 50.33 50.00 53.78 SEED-Bench 60.44 60.28 59.31 57.13 46.60 46.45 49.09 SEED-Bench (Video) 39.02 41.90 38.86 41.02 32.50 33.01 31.33
DiT SD3 SD2.1 SigLIP C+D@224 C+D@336 MMBench 33.68 32.82 28.87 61.86 65.72 65.12 MME 902.00 843.43 905.27 1425.00 1436.42 1475.19 MMMU 32.7 32.4 32.8 35.8 36.9 34.6 OKVQA 33.75 34.95 34.41 54.01 55.94 56.92 TextVQA 10.82 10.77 10.46 36.00 40.04 46.17 VizWiz 49.92 47.12 46.59 53.17 54.04 53.44 ScienceQA (Full) 66.62 66.14 65.69 69.11 70.87 69.45 ScienceQA 68.12 67.98 67.13 70.17 71.70 70.31 ScienceQA (Img) 63.46 62.27 62.67 66.88 69.11 67.63 SEED-Bench (Img) 40.66 38.94 38.82 64.40 65.39 66.38 SEED-Bench 38.31 36.96 36.84 59.41 60.47 61.39 SEED-Bench (Video) 29.41 29.46 29.33 40.48 41.84 42.48
Table 4: Benchmark performance of all 13 settings. C+D means feature combination of CLIP and DINOv2.
@@ A.4 All Settings AC Scores
We provide the AC scores of all 13 vision representation settings, as summarized in Table 5.
CLIP@336 CLIP@224 OpenCLIP DINOv2 SDim SD1.5 SDXL Correspondence PCK@0.10 15.66 14.3 16.22 24.51 20.9 22.02 16.52 Cross-modal Alignment MMBench 0.788 0.815 0.524 0.462 0.349 0.351 0.357 MME 0.791 0.818 0.532 0.472 0.370 0.366 0.358 MMMU 0.782 0.813 0.512 0.464 0.356 0.353 0.358 OKVQA 0.801 0.825 0.543 0.472 0.379 0.375 0.353 TextVQA 0.797 0.820 0.536 0.458 0.374 0.367 0.347 VizWiz 0.794 0.815 0.530 0.469 0.370 0.372 0.360 ScienceQA (Img) 0.799 0.826 0.541 0.472 0.371 0.356 0.342 SEED-Bench (Img) 0.810 0.829 0.554 0.491 0.376 0.359 0.342
DiT SD3 SD2.1 SigLIP C+D@224 C+D@336 Correspondence PCK@0.10 1.91 3.09 6.99 12.89 23.62 26.08 Cross-modal Alignment MMBench 0.387 0.374 0.348 0.505 0.537 0.512 MME 0.392 0.377 0.338 0.526 0.536 0.511 MMMU 0.398 0.363 0.333 0.499 0.526 0.504 OKVQA 0.366 0.390 0.351 0.540 0.537 0.514 TextVQA 0.368 0.387 0.347 0.532 0.530 0.506 VizWiz 0.383 0.401 0.348 0.527 0.525 0.505 ScienceQA (Img) 0.363 0.383 0.358 0.541 0.537 0.512 SEED-Bench (Img) 0.367 0.388 0.371 0.549 0.545 0.525
Table 5: AC scores of all 13 settings. C+D means feature combination of CLIP and DINOv2.
@@ A.5 More Visualization of Correspondence
We provide additional visualizations of correspondence for four different vision representations: CLIP, SigLIP, DINOv2, and Stable Diffusion 1.5. Figures 6 and 7display pairs of source-target images for each of the four vision representations. In each pair, the left image is the source, and the right image is the target. The red dot on both images indicates the predicted key points using the vision representation. Ideally, these key points should correspond to the same semantic meaning. For example, a red dot on the ”left cat ear” in the source image should correspond to the ”left cat ear” in the target image. The green areas highlight regions of relatively high similarity with the source points.

### v2

@@ Results.
In Figure 4, we demonstrate that the AC policy consistently predicts the optimal vision representation using minimal resources within a finite search space of 15 settings. Our aim is to finetune only a small subset of this space while ensuring that the optimal vision representation is among the top-3 predictions (Recall@3). With a computational budget equivalent to 4 full finetuning runs, a random subset selection achieves only 26.7% Recall@3. In contrast, the AC policy achieves 87.72% Recall@3 (averaged over 6 benchmarks) while still requiring just 4 full training runs. For further details, see Section 5.5.
Figure 4: Given a limited budget of 4 finetunings, AC policy achieves 87.72% Recall@3 in predicting the optimal vision representation.
@@
SigLIP ViT-B/16  (Zhai et al., 2023b) 224
SigLIP ViT-L/16  (Zhai et al., 2023b) 256
SigLIP2 ViT-L/16  (Tschannen et al., 2025) 256
@@
[…]
tableVision representations explored.
5 Empirical Result Details
5.1 Experiment Settings
For our MLLM pipeline, we deploy a LLaMA-based LLM, specifically Vicuna-7B 1.5 (Zheng et al., 2023), and utilize a widely used 2-layer GeLU-MLP connector as the projector. For vision representation, we explore a variety of encoder types and sizes, input resolutions, training paradigms, and feature combinations, as detailed in Table 4.
Our training process consists of two stages. In Stage 1, we perform alignment using the LLaVA 1.5 dataset with 558K samples (Liu et al., 2024e), training only the projector. In Stage 2, we train both the connector and the language model on an expanded LLaVA 1.5 dataset containing 665K samples.
The MLLM benchmarks used in this paper include four vision-based benchmarks (MMBench (Liu et al., 2023), MME (Fu et al., 2023), OKVQA (Marino et al., 2019), SEED-Bench (Li et al., 2024b)) and four QCR-based benchmarks (MMMU (Yue et al., 2024), TextVQA (Singh et al., 2019), VizWiz (Gurari et al., 2018), ScienceQA (Lu et al., 2022)).
@@ 5.2 AC Score
To compute the cross-modal alignment score, we perform Stage 1 training on all vision representations to obtain the projected vision representations. This stage requires significantly less computation than Stage 2, involving only 0.298% of the trainable parameters. The image–caption pairs are taken from the LLaVA-558K dataset, and the alignment score is computed by averaging the results across 100 randomly sampled images.
For the correspondence score, we follow common practices using the SPair-71k dataset (Min et al., 2019). Note that all benchmarks share the same A and C scores, while the quadratic parameters in the function fitting adapt to capture variations across tasks.
@@ 5.3 Feature Extraction
Diffusion model is primarily used for generating images via multi-step denoising, yet a recent trend is to use diffusion model as the vision representation model (Xu et al., 2024; 2023; Zhang et al., 2024c; Tong et al., 2024a). Specifically, for diffusion models, given an image $I\in\mathbb{R}^{H\times W\times 3}$, we first add noise to the VAE-encoded representation of $I$:
@@ 5.4 Additional Results on the Law of Vision Representation
Random 4.03% 1.75%
A Score 80.53% 58.00%
C Score 39.02% 14.57%
AC Score 80.55% 62.06%
@@
Random 36.90% 31.26%
A Score 91.48% 79.67%
C Score 56.56% 30.11%
AC Score 94.06% 83.85%
@@
In Section 3, we demonstrate the strong correlation between the AC score and MLLM performance by analyzing the coefficient of determination ($R^{2}$) obtained from fitting a quadratic regression model. In this section, we further ablate the experiments by adding baselines, fitting model performance with random scores, A scores, and C scores separately. Additionally, we explored the relationship between the A score and C score by fitting a linear regression model. We avoid higher-degree transformations to prevent overfitting, which could obscure the true relationship between A and C scores.
As shown in Table 1, the results indicate that using the AC score consistently outperforms all other settings in terms of $R^{2}$ values. While this observation holds regardless of the degree of fitted function, using a second-degree polynomial on A and C scores yields the highest correlation with model performance. This suggests an inherent trade-off between A and C scores: vision representations with high cross-modal alignment often exhibit lower correspondence, and vice versa.
Interestingly, we observe a lower correlation between OCR-based benchmark performance and C scores, which leads to a reduced correlation between the AC score and OCR-based benchmark performance. In Section 6, we discuss how the use of the SPair-71k correspondence dataset across all benchmarks fails to adequately capture correspondence in images containing text.
@@ 5.5 Additional Results on the AC Policy
In Section 4, we demonstrate that fitting the AC score consistently predicts the optimal vision representation with minimal resources, given a finite search space—in this case, 15 settings. In this section, we provide detailed visualization for Figure 4.
@@
When performing ablation experiments on vision encoders, it’s common to randomly select a subset to train on. However, as shown in Figure 5, with 1000 runs of simulated ablation experiments, we found that to include the optimal vision representation 85.6% of the time, at least 13 out of the 15 settings need to be trained. This suggests that running a small subset of vision representations is unreliable, especially as the search space expands, making it increasingly unlikely to identify the true optimal representation by training only a subset.
@@
In contrast, the AC policy requires only 4 full training runs on average to reach 87.72% Recall@3. For the most successful prediction benchmark, OKVQA, the policy successfully identifies the optimal configuration among the top three choices in 91.7% of cases, with only four language model finetuning runs across a 15-setting search space. This result shows that AC policy significantly reduces the effort and cost of exploring vision representations for MLLMs.
6 Limitation
Figure 6: Visualization of correspondence on natural images and images containing text for CLIP and DINOv2.
We find that OCR‐based benchmarks correlate less with the AC score than vision‐based ones, making MME and MMMU outliers. For example, SigLIP2‐Large@256 outperforms CLIP‐Large@336 in both alignment (-1.81 vs. -1.97) and correspondence (16.75 vs. 15.66) but underperforms on MME due to OCR‐heavy categories.
This discrepancy arises because vision representations exhibit different correspondence accuracy on different domain of images, as shown in Figure 6. The SPair‐71k dataset (for the C score) focuses on natural images (e.g., cats, trains), whereas CLIP excels at text‐based tasks not captured in SPair‐71k. Likewise, our alignment score, calculated using LLaVA 558K (a subset of LAION, CC, SBU), lacks sufficient OCR, numeric, and symbolic data. Consequently, both the A and C scores underrepresent performance on OCR‐related tasks.
To our knowledge, an OCR-specific correspondence dataset does not currently exist, and systematically designed OCR short caption datasets are scarce. We intend to pursue further investigation in this direction and encourage other researchers to do the same, as advancements in this area would be valuable for the broader MLLM community—particularly for tasks requiring the understanding of tables and charts, a fundamental capability.
@@ A.1 Theoretical Justification of Vision Representation with High Cross-modal Alignment
Since our language model $f$ is well-trained and pre-normed, the input space to each transformer layer is bounded and compact, meaning that the values of the input are bounded by a small constant. This implies that the continuously differentiable function $f$ is Lipschitz (Kim et al., 2021). This property ensures that small changes in the input of the language model of the MLLM result in small, controlled changes in the output:
$\Big{\|}f\big{(}[E_{c}^{\text{image}},E_{1},E_{2},\dots,E_{n}]\big{)}$ $-f\big{(}[E_{c}^{\text{text}},E_{1},E_{2},\dots,E_{n}]\big{)}\Big{\|}$
$\leq L\Big{\|}[E_{c}^{\text{image}},E_{1},E_{2},\dots,E_{n}]-[E_{c}^{\text{text}},E_{1},E_{2},\dots,E_{n}]\Big{\|}$
$\leq L\epsilon.$ (7)
@@ A.3 All Settings Benchmark Performance
In this section, we present the performance results of all 15 vision representation settings, as summarized in Table 2. The benchmarks we evaluated include:
@@
  • SEED-Bench (Li et al., 2024b): A benchmark consisting of multiple-choice questions designed to assess both spatial and temporal understanding.
CLIP@336 CLIP@224 OpenCLIP SigLIP Base SigLIP Large MMBench 64.26 64.18 63.40 61.86 65.46 MME 1502.70 1449.64 1460.28 1425.00 1455.18 MMMU 35.0 36.2 37.2 35.8 36.6 OKVQA 53.20 56.13 56.36 54.01 57.02 TextVQA 46.04 42.67 40.13 36.00 42.44 VizWiz 54.27 51.69 52.11 53.17 51.49 ScienceQA 69.26 68.82 67.87 66.88 70.20 SEED-Bench 66.09 65.13 64.71 64.40 66.28
C+D@224 C+D@336 SigLIP2 Large DINOv2 DiT MMBench 65.72 65.12 67.35 58.50 33.68 MME 1436.42 1475.19 1486.66 1295.47 902.00 MMMU 36.9 34.6 34.2 34.6 32.7 OKVQA 55.94 56.92 57.57 54.78 33.75 TextVQA 40.04 46.17 47.2 14.27 10.82 VizWiz 54.04 53.44 56.55 49.67 49.92 ScienceQA 69.11 67.63 69.41 65.15 63.46 SEED-Bench 65.39 66.38 68.22 61.39 40.66
SDXL SD3 SD2.1 SD1.5 SDim MMBench 43.73 32.82 28.87 42.53 52.84 MME 1212.69 843.43 905.27 1163.90 1205.33 MMMU 32.8 32.4 32.8 33.9 33.7 OKVQA 41.78 34.95 34.41 39.14 46.04 TextVQA 11.81 10.77 10.46 11.64 13.77 VizWiz 47.14 47.12 46.59 50.14 47.33 ScienceQA 65.25 62.27 62.67 63.31 66.34 SEED-Bench 53.78 38.94 38.82 50.00 50.33
Table 2: Benchmark performance of all 15 settings. C+D means feature combination of CLIP and DINOv2. The table provides data points for function fitting and is not intended for comparison.
@@ A.4 All Settings AC Scores
We provide the AC scores of all 15 vision representation settings, as summarized in Table 3.
CLIP@336 CLIP@224 OpenCLIP SigLIP-B SigLIP-L Correspondence PCK@0.10 15.66 14.30 16.22 12.89 13.66 Cross-modal Alignment Log likelihood -1.97 -1.98 -1.93 -1.92 -1.83
C+D@224 C+D@336 SigLIP2-L DINOv2 DiT Correspondence PCK@0.10 23.62 26.08 16.75 24.51 1.91 Cross-modal Alignment Log likelihood -1.96 -1.95 -1.81 -2.32 -3.76
SDXL SD3 SD2.1 SD1.5 SDim
Correspondence
PCK@0.10 16.52 3.09 6.99 22.02 20.90
Cross-modal Alignment
Log likelihood -2.69 -4.13 -2.81 -2.53 -2.37
Table 3: AC scores of all 15 settings. C+D means feature combination of CLIP and DINOv2. The table provides data points for function fitting and is not intended for comparison.
@@ A.5 More Visualization of Correspondence
We provide additional visualizations of correspondence for four different vision representations: CLIP, SigLIP, DINOv2, and Stable Diffusion 1.5. Figures 7 and 8 display pairs of source-target images for each of the four vision representations. In each pair, the left image is the source, and the right image is the target. The red dot on both images indicates the predicted key points using the vision representation. Ideally, these key points should correspond to the same semantic meaning. For example, a red dot on the ”left cat ear” in the source image should correspond to the ”left cat ear” in the target image. The green areas highlight regions of relatively high similarity with the source points.
@@
A.6 Pseudo Code
Computing the A score is a simple loss calculation using the frozen MLLM; therefore, we provide pseudocode for the other algorithms, including the computation of the C score at Algorithm 1, region-based sampling at Algorithm 2, and the AC policy at Algorithm 3.
Input: Set of paired images with key points $S$ from SPair-71k;
Vision encoder $E$;
Threshold $T$.
Output: $C$ score for vision encoder $E$.
// Initialize correspondence lists
$G\leftarrow[]$ ;
  // Ground truth keypoint correspondences
$P\leftarrow[]$ ;
  // Predicted keypoint correspondences
foreach $(I_{1},K_{1},I_{2},K_{2})\in S$ do
       // Extract feature representations
       $F_{1}\leftarrow E(I_{1})$;
       $F_{2}\leftarrow E(I_{2})$;
      // Compute similarity matrix
       $S_{\text{sim}}\leftarrow F_{1}\cdot F_{2}^{T}$;
      // Transform keypoints from $I_{1}$ to $I_{2}$
       $\hat{K}_{2}\leftarrow\text{calculate\_keypoint\_transformation}(S_{\text{sim}},K_{1})$;
      // Store ground truth and predicted keypoints
       $G.\text{append}(K_{2})$;
       $P.\text{append}(\hat{K}_{2})$;
// Compute correctness score
$E_{\text{error}}\leftarrow\text{Euclidean\_distance}(P,G)$;
$C_{\text{correct}}\leftarrow\text{sum}(E_{\text{error}}<T)$;
$C_{\text{score}}\leftarrow\frac{C_{\text{correct}}}{\text{total keypoints in }K_{2}}$;
Algorithm 1 Compute C Score
Input: $k$ A and C score pairs from models $ACs$; $past\_sampled$ models; current sampling $level$ ($1$ to $k^{\prime}$, increments when regions are exhausted as each region is sampled only once)
Output: Sampled $model$ to train next
$regions\leftarrow\{\}$;
for $AC\in ACs$ do
       $region\_key\leftarrow\text{determine\_region}(A,C,level)$ ;
        // Identify the region based on A and C coordinates
       $regions[region\_key].append((model,A,C))$;
Remove models in $past\_sampled$ from $regions$;
$remaining\_regions\leftarrow\text{keys of }regions$;
$chosen\_region\leftarrow\text{randomly select from }remaining\_regions$;
$model\leftarrow\text{randomly select from }regions[chosen\_region]$;
Algorithm 2 Region-based Sampling
Input: $k$ vision encoders with pretrained projectors $V$; computation budget $k^{\prime}$
Output: A ranking of $k$ MLLMs based on performance
$ACs\leftarrow[(\textnormal{{Compute\_A\_Score}}(v),\textnormal{{Compute\_C\_Score}}(v))\mid v\in V]$;
$past\_sampled\leftarrow[]$;
$train\_ACs\leftarrow[]$;
$train\_performance\leftarrow[]$;
for $i\leftarrow 1$ to $k^{\prime}$ do
       $model\leftarrow\textnormal{{Region\_based\_Sampling}}(ACs,past\_sampled)$;
       $performance\leftarrow\text{Fully train }model$;
       $train\_ACs.\text{append}(\text{AC of }model)$;
       $train\_performance.\text{append}(performance)$;
       $past\_sampled.\text{append}(model)$;
$poly\leftarrow\text{PolynomialFeatures}(degree=2)$;
$transformed\_train\_ACs\leftarrow poly.\text{fit\_transform}(train\_ACs)$;
[…]
$regression.\text{fit}(transformed\_train\_ACs,train\_performance)$;
$ranking\leftarrow\text{Rank }V\text{ by regression predictions on }ACs$;
Algorithm 3 AC Policy
A.7 Limitation of AC Policy
Figure 9 shows two benchmarks—MME and MMMU—where the AC policy fails to predict the optimal vision representation. For details on the reason of this behavior, please refer to Section 6.
Figure 9: AC policy.
A.8 AC Policy Recall@1
In Section 5.5, we used Recall@3 as a metric for AC policy effectiveness to capture potential performance fluctuations during MLLM training and evaluation. Here, we also report Recall@1 for predicting the optimal vision representation. As shown in Figure 10, AC policy consistently outperforms random testing even under this stricter metric.
Figure 10: Number of full training (LLM finetuning) cycles required to include the optimal vision representation within the top-1 predictions (Recall@1).